# Securing Automated Insulin Delivery Systems: A Review of Security Threats and Protectives Strategies

### v1

@@ 3.1.4 Subcutaneous Insulin Pump
A subcutaneous insulin pump includes a reservoir, which stores insulin and supplies it to the delivery mechanism, an infusion set consisting of a thin tube and a subcutaneous cannula for insulin delivery, and a microcontroller unit that orchestrates the pump’s functions by executing programmed commands based on input from the control platform. Security concerns associated with the pump arise from device malfunctions (e.g., infusion site failures)[43][44] and malicious control I/O operations of the pump[13].
3.2 Regulation Standards
Applied Area Regulation Standards Cybersecurity Note US Policy for Device Software Functions and Mobile Medical Applications Require adequate cybersecurity to secure the correct functioning of medical devices Provide specific cybersecurity considerations Section 3060 and 3060(a) of the 21st Century Cure Acts EO 14019 on America’s Supply Chains Require detailed technical cybersecurity measures used by the product and related services Not mandated for medical devices yet; could be in the future Guidance for FDA staff Cybersecurity in Medical Devices Provides expectations for inspection and approval, but no detailed federal guidance EU Medical Devices Regulation (MDR) No provisions specifically labeled ‘cybersecurity’ Highest priority MDR Annex1: General Safety and Performance Requirements (GSPR) Require that medical devices do not fail due to any type of adversary attacking them Imply minimum levels of encryption, access control, and physical cybersecurity Network and Information Systems (NIS 2) Directive Cybersecurity requirement of general critical IT infrastructure Integrated into Member States in various ways Cyber Resilience Act and the Cyber Solidarity Act Regulate the cybersecurity of any network-enabled devices Supposed to entail more resilience (greater recovery from failure) than MDR Security standards: ISO 14155:2011 Mostly developed by private actors and does not necessarily reflect best practices. Allow usage of the standard for a fee, which becomes a barrier to adoption General Data Protection Regulation (GDPR) Protecting the fundamental right to privacy Only considering cybersecurity requirements related to personal data
Table 3: Regulation standards of cybersecurity in medical devices.
@@ 3.2 Regulation Standards
The regulatory landscape governing AID systems cybersecurity is evolving, with growing recognition of safeguarding patient safety. Table 3 summarizes the core regulatory standards for AID systems cybersecurity in the United States (US) and European Union (EU). Given the global nature of medical device markets, the regulatory frameworks of many other countries closely align with these standards.
In the US, the FDA is responsible for overseeing medical devices’ compliance with regulations. Existing referable legal sources and guidances for FDA staff underscore the importance of cybersecurity as a key criterion in the FDA’s device inspection and approval processes [45]. In the Policy for Device Software Functions and Mobile Medical Applications, medical device manufacturers are required to implement adequate measures to protect against potential cyberattacks. Nonetheless, while the emphasis on cybersecurity is clear, there remains a notable absence of detailed, technical mandates similar to those applied in supply chain regulations. Given the increasing focus on medical device security, more specific technical requirements may be introduced in the future [46].
In the EU, the Medical Devices Regulation (MDR) serves as the main legislative framework, along with other cybersecurity standards. Although the MDR is positioned as the highest priority, it contains no provisions specifically labelled ‘cybersecurity’. Nevertheless, Annex 1 of the MDR, General Safety and Performance Requirements (GSPR) and other cybersecurity legislation provide a promising landscape regulation for cybersecurity in medical devices. For example, GSPR requires medical devices not to fail due to any type of security attack and that the necessary cybersecurity for functioning must be in place.
Generally, the EU and the US regulations mainly provide high-level principles rather than detailed technical specifications. While manufacturers are encouraged to adopt best practices like encryption, data access control, and software patching, there is no mandated standard for real-time monitoring or automatic threat response systems. A more specific review of cybersecurity in medical device regulation and its future consideration can be found in [46].
Beyond government regulations, a nonprofit organization named Diabetes Technology Society (DTS) proposed a Cybersecurity Standard for the Connected Diabetes Devices project [47]. This standard provides a standardized approach to thoroughly evaluating and confirming the conformity and confidence that the medical device can protect itself against applicable security threats, offering an example of how regulation could evolve toward more comprehensive and concrete evaluations.
3.3 Commercial Products
Within the regulation of existing cybersecurity frameworks, leading commercial AID system products demonstrate only basic cybersecurity capabilities, lacking comprehensive and robust protective measures, as outlined in Table 4. The DBLG1 System employed data encryption to secure wireless communication and Medtronic 780G includes a safety feature that pauses insulin delivery when the glucose level falls below a preset threshold, while others do not specify particular cybersecurity strategies.
Experiments [12][48] have demonstrated that at least two commercial products lack standard cryptographic mechanisms and the wireless communication protocol can be fully reverse-engineered, allowing eavesdropping, replay and command manipulation attacks to be launched. These vulnerabilities have led to commercial products being recalled [10] and warning of use [49]. Even devices with basic encryption remain susceptible to attacks that compromise integrity and availability. In the meantime, mechanisms hired to prevent patient harm are insufficient. Process-aware attacks [40] and personalized insulin dose manipulation attacks [50] can cause severe health consequences by subtly altering insulin delivery without injecting over-dose insulin to raise alarms.
Product Components Main Functions Cybersecurity CamAPS FX User’s smartphone (controller);
DANA/Ypsomed insulin pump;
Dexcom/Libre 3 CGM
Automated basal insulin delivery; Manual bolusing for meals User’s responsibility
Medtronic 780G MiniMed Insulin pump (controller integrated);
MiniMed CGM
Automated basal insulin delivery; Recommend precise bolus doses Mainly the user’s responsibility; Pause delivery when the glucose level is less than a preset level
Control IQ User’s smartphone (controller);
Tandem t: Slim insulin pump;
Dexcom G6/G7 CGM;
Automated basal insulin delivery; Automatic correction boluses delivery User’s responsibility
Omnipod 5 User’s smartphone (controller);
Omnipod insulin pumps;
Dexcom G6 CGM
Automatic basal insulin adjustments User’s responsibility
DBLG1 System DBLG1 controller;
Kaleido pump;
Dexcom G6 CGM
Automated basal insulin delivery; Automatic correction boluses delivery Mainly the user’s responsibility; Data encrypted
Table 4: Cybersecurity in representative AID system products.
Currently, according to product manuals, the responsibility for detecting and responding to security threats is placed on the patients themselves. However, this expectation is often impractical in real-world scenarios, particularly during sleep or daily activities. Additionally, the absence of a CGM display further limits the user’s ability to promptly detect deviations in BG readings from the control platform. Under these circumstances, AID system failures remain prevalent. A thesis [25] listed 24 documented device recalls and failures associated with a single AID systems manufacturer up until 2023 (Table 05). Despite evidence of their effectiveness in diabetes management, a cross-sectional survey [51] indicates that concerns over privacy and reliability may hinder AID systems adoption.
Given these challenges, it is imperative for both manufacturers and regulators to establish comprehensive and specific cybersecurity standards. Ensuring that AID systems devices can timely detect and mitigate security threats — without relying solely on user intervention—will be essential for enhancing system reliability, fostering user trust, and fully realizing the potential of AID systems in diabetes care.
4 Attack Vectors in AID systems
@@
Building on the discussion of cybersecurity vulnerabilities in AID systems, it is crucial to examine the specific cyber threats that target these systems. Table 5 presents a chronological overview of both real-world attacks and emerging threats identified in the literature on AID systems.
@@
As illustrated in the taxonomy (Figure 2), malicious attacks on AID systems can be broadly categorized into cyber-physical attacks and process-aware attacks, based on whether they exploit knowledge of the system’s dynamic control processes [52]. Cyber-physical attacks, which are common across various IoMT devices, compromise system functionality by targeting software, hardware, or communication channels without requiring an in-depth understanding of the underlying control mechanisms. In contrast, process-aware attacks specifically target semi- or closed-loop medical systems such as AID. By making subtle manipulations within the operational bounds of the system, these attacks can significantly impact patient safety [40]. Since cyber-physical attacks align with broader IoMT security concerns, threats and potential countermeasures have been detailed in existing surveys [29][30][31], this section focuses on analysing attack vectors and technical details that are relevant to AID systems.
To systematically analyse these attack vectors, we categorize them into three key areas based on their impacts on the fundamental principles of cybersecurity: confidentiality, integrity, and availability. Table 6 presents an overview of these attack vectors, along with brief descriptions and feasible countermeasures.
Attack Name Ref. Type Cause of Failure Impact Adversarial Capabilities RL-based Control Model Attack Chang et al. [53] (10/2024) Literature (Experimentally Validated) The control system is vulnerable to anomalous inputs. A maximum CGM reading of 400mg/dL caused high rates (over 40%) of hypoglycaemic failure within a few hours; a very slight perturbation (±1mg/dL) can lead to disproportionate patient harm over time High - Read access to the weights of the trained model or estimated weights AI/ML model attacks Elnawawy et al. [54] (01/2024) Literature (Experimentally Validated) Introducing adversarial CGM data points during inference through a known vulnerability in the Bluetooth communication channel Wrong diagnosis or treatment Medium - Can exploit the Bluetooth communication stack to intercept and manipulate the CGM measurements Unauthorized Remote Control Medtronic [10] (01/2023) Real-world (Recall) The settings and insulin delivery of the pump could be potentially controlled through a wireless connection nearby Overdosing or Underdosing Medium - Intercept and manipulate the communication between devices Process-Aware Attacks Stergiopoulos et al. [40] (11/2023) Literature (Experimentally Validated) Utilising knowledge of the dynamic process to alter sensory data or the logic inside the controller in an intended but subtle way Overdosing or Underdosing High - Prior knowledge of the components’ implementation mechanism Replay Attacks Walenda et al. [55] (11/2023) Literature (Experimentally Validated) Sending repeating insulin dosages to the insulin pump through a vulnerability in the wireless protocol Replay attacks have the greatest effect when the injected insulin doses are high Low – Intercept the wireless connection and report past packets DoS Attacks Walenda et al. [55] (11/2023) Literature (Experimentally Validated) Overload the system to stop insulin delivery for extended periods DoS attacks are most effective at low, increasing doses, causing longer out-of-range times and greater spikes and dips than replay attacks. Low - Jam the communication channel Personalized insulin dose manipulation attack Levy-Loboda et al. [11] (08/2022) Literature (Theoretical) Craft attacks that only threat target patients and are capable of evading extreme overdose/underdose detection Overdosing or Underdosing High – Prior knowledge and manipulation of the insulin dosage Bias Injection Attacks Emre Tosunal [50] (06/2022) Literature (Theoretical) A constant bias is injected into the compromised sensor readings Overdosing or Underdosing Medium - Intercept and manipulate the communication between devices Ripple False Data Injection Attacks Kashyap et al. [56] (05/2021) Literature (Experimentally Validated) Using minimal input perturbations to stealthily change the DNN-predicted insulin dosage More allowed perturbations make attacks easier to find for targeted attacks, with discovery time as short as 60ms. Random attacks can easily induce 10–20 mg/dL output changes across scenarios. High - Read access to the weights of the trained model or estimated weights Command Manipulation Attack Khan et al. [41] (12/2020) Literature (Theoretical) Change the instructions or data that implement the algorithmic steps Overdosing or Underdosing High - Access to the relevant resources; modify a command (i.e., instructions that implement the algorithmic steps) Unauthorized Remote Control Cooke et al. [13] (04/2020) Literature (Experimentally Validated) An adversary can remotely control insulin delivery and access unencrypted medical data Overdosing or Underdosing Medium - Intercept and manipulate the communication between devices Replay Attacks Medtronic [49] (08/2018) Real-world (Notification) Involuntary bolus insulin could be delivered to the pump user by copying the wireless radio frequency (RF) signals from remote controllers in proximity Overdosing Low – Intercept the wireless connection and report past packets CGM Attacks Reverberi and Oswald [57] (08/2017) Literature (Experimentally Validated) The Dexcom G4 protocol does not employ cryptographic algorithms Enables the tracking of a user, the jamming of the communication channels and the forging of incorrect sensor readings Low to medium - software-defined radio and low-cost RF chipsets, reverse engineering Saturation-based Sensor Spoofing Attacks Park et al. [58] (08/2016) Literature (Experimentally Validated) Make a sensor ignore legitimate inputs by using an additional infrared source to saturate the sensors Inject up to 3.33 times the intended amount of fluid or 0.65 times for a 10-minute period Low – Saturating sensor from an external source Software radio-based attacks Marin et al [48] (03/2016) Literature (Experimentally Validated) Fully reverse-engineered the wireless communication protocol between all the peripherals of the insulin pump system Compromise both the security and privacy of the patient Passive attacks: Low - Radio interception and generation device
@@
Software radio-based attacks Li et al [12] (07/2011) Literature (Experimentally Validated) No standard cryptographic mechanisms; the wireless protocol can be fully reverse-engineered, enabling eavesdropping, replay, and command manipulation attacks. Compromise both the security and privacy of the patient Passive attacks: Low - Radio interception and generation device
@@
Table 5: Documented attacks on AID systems.
Security Aspect Attack Vector Description Countermeasures Confiden-tiality Eavesdropping
@@
Proximity communication channel;
Secure protocols;
Authentication
Do-it-yourself hacking[16] Patients extract data not automatically provided by products’ software Authentication Integrity Bias injection attack[39][50] A constant bias was injected into the sensor readings Intrusion detection systems;
@@
Infusion patterns mining;
Control strategies assess
False data injection attack[44][10] Change transmitted data with random or synthesized data Replay attacks[6] Intercept and replay legitimate data AI/ML model attacks[54] Deviate the ML model’s prediction during training or inference Computational code attacks[41] Compromise the command computational code Formal specification Malicious pump driver [59] Change pump I/O operations Avail-ability Denial-of-Service (DoS) /Distributed DoS [59] Flooding the target with traffic results in the inability to respond timely Authentication;
Secure protocols;

### v2

@@ III-A4 Subcutaneous Insulin Pump
A subcutaneous insulin pump includes a reservoir, which stores insulin and supplies it to the delivery system, an infusion set consisting of a thin tube and a subcutaneous cannula for insulin delivery, and a microcontroller unit that orchestrates the pump’s functions by executing programmed commands based on input from the control platform. Security concerns associated with the pump arise from device malfunctions (e.g., infusion site failures)[88][86] and malicious control I/O operations of the pump[27].
III-B Regulation Standards and Guidelines
@@
The regulatory landscape governing the cybersecurity of AID systems is evolving, with growing recognition of cybersecurity as a critical component of patient safety. This section examines the regulatory frameworks of the United States (US) and the European Union (EU). Given the global nature of medical device markets, many other countries have adopted regulations that closely align with these standards. Appendix A summarises the core regulatory standards and guidelines relevant to AID system cybersecurity in these regions.
In the US, the FDA oversees regulatory compliance for medical devices. Existing legislation and FDA guidance documents highlight cybersecurity as a key consideration in approval and inspection processes. Notably, the FDA has issued guidance for premarket submissions titled ”Cybersecurity in Medical Devices: Quality System Considerations and Content of Premarket Submissions” [38], and for postmarket submissions titled ”Postmarket Management of Cybersecurity in Medical Devices” [132]. These documents provide prescriptive and structured approaches for manufacturers to conduct risk management and implement cybersecurity controls. In addition, the NIST’s cybersecurity framework [95] provides a structured catalogue of typical cybersecurity activities and controls that enterprises should consider.
In the EU, the Medical Devices Regulation (MDR) [83] serves as the primary legislative framework, supplemented by other cybersecurity standards. Annexe I of the MDR, General Safety and Performance Requirements (GSPR), requires that medical devices must not fail due to security attacks and that necessary cybersecurity protections must be in place. To facilitate implementation, the Medical Device Coordination Group (MDCG) issued the guidance document “MDCG 201916: Guidance on Cybersecurity for medical devices” [41], which provides recommendations to manufacturers on fulfilling MDR cybersecurity requirements. Compared with FDA guidance, MDCG offers relatively high-level principles on how to operationalise these activities [120].
While cybersecurity regulations in the US and EU both require manufacturers to identify, assess, and mitigate cybersecurity risks across the product lifecycle, responsibility for comprehensive threat modelling and protective measures implementation rests with the manufacturer. So the approach and depth are determined by the manufacturer and subject to regulatory review. This creates challenges in mapping the broad landscape of security requirements to specific devices and maintaining up-to-date threat models. Moreover, current guidelines lack formal mechanisms for establishing confidence in a product or system’s capacity to defend against or recover from cyber attacks through independent assessment [65].
To address these limitations, independent organisations and researchers have proposed complementary approaches. For example, the nonprofit organisation Diabetes Technology Society (DTS) introduced the DTSec (DTS Cybersecurity Standard for Connected Diabetes Devices) to define AID-specific security requirements and enable independent compliance assessment [32]. In addition, to assess the cyber risk against the up-to-date threat landscape, Shen [117] proposed a cyber risk assessment method that evaluates Manufacturers Disclosure Statement for Medical Device Security (MDS2) data sheets against updated cyber threat intelligence using MITRE ATT&CK. These initiatives highlight potential directions for evolving regulation toward more comprehensive and concrete cybersecurity compliance regulations.
III-C Commercial Products
With the rising prevalence of diabetes and continued technological advancement, global demand for AID systems is rapidly increasing, with the market valued at US\$2.34 billion in 2023 and projected to reach US\$5.80 billion by 2032 [50]. Nevertheless, safety and security challenges persist. For example, by 2023, 24 device failures and recalls were documented for a single AID system manufacturer [59]. A cross-sectional survey [6] revealed that concerns regarding privacy and reliability are one of the major factors that hinder broader adoption of AID systems. Appendix B summarises the core components, functions, and security measures of representative products. Although the cybersecurity capabilities of leading products have improved, they remain limited in terms of security protection, risk management, and monitoring.
Experimental studies have demonstrated that two early-generation AID products lack standard cryptographic mechanisms, allowing full reverse-engineering of their wireless protocols and exposing them to eavesdropping, replay, and data manipulation attacks[73][81]. These vulnerabilities resulted in official warnings of use [85] and product recalls [84]. Although more recent products employ encryption and authentication to protect wireless communication, the level of security depends on factors such as method strength, key management, and system integration. Even with strong encryption and authentication, vulnerabilities remain to advanced threats like side-channel attacks, denial-of-service (DoS) attacks, and physical tampering.
To address advanced or unknown threats that may bypass safeguards, robust risk management is essential for ensuring device reliability under attack. However, risk management functions in current products are limited. Insulin pumps incorporate preset minimum and maximum dosage and infusion rate limits, while systems like the Medtronic 780G suspend insulin delivery when glucose levels drop below a predefined threshold. Yet, these threshold-based approaches are ineffective against more sophisticated attacks. Research has demonstrated that personalised insulin dose manipulation attacks can subtly alter insulin delivery in ways that are harmful to specific patients at particular times while remaining within general safety constraints [72]. In addition, false data injection attacks can falsify glucose readings to suppress alarm activation, and computational attacks can modify algorithm logic [27]. Detecting these attacks requires identifying abnormal patterns rather than comparing with predefined thresholds.
Meanwhile, the main responsibility for monitoring and securing devices is explicitly assigned to patients themselves as indicated in user manuals [22]. However, this expectation is often unrealistic, especially during sleep or other daily activities. Besides, the absence of real-time displays in CGM systems further limits patients’ ability to detect discrepancies between glucose readings and data relayed to the control platform.
Given these limitations, there is a pressing need for manufacturers to establish sufficient security measures to prevent, detect, and respond to security threats in real time without relying solely on patient intervention. Strengthening security in this manner is essential to improving system reliability and fostering user trust.
IV Attack Vectors in AID systems
@@
Building on the discussion of security vulnerabilities in AID systems, it is crucial to examine the specific attack vectors that target these systems. Table III presents a reverse chronological overview of real-world attacks and emerging threats identified in the literature, detailing their causes, impacts on device functionality, and the adversarial capabilities required.
@@
As illustrated in the taxonomy (Figure 2), malicious attacks on AID systems can be broadly categorised into cyber-physical attacks and process-aware attacks, based on whether they exploit knowledge of the system’s dynamic control processes [63]. Cyber-physical attacks, which are common across various IoMT devices, compromise system functionality by targeting software, hardware, or communication channels without requiring an in-depth understanding of the underlying control mechanisms. In contrast, process-aware attacks specifically target semi- or closed-loop control like AID systems. By making subtle manipulations within the operational bounds, these attacks can significantly impact system behaviours to threat patient safety [121]. Since cyber-physical attacks align with broader IoMT security concerns, threats and potential countermeasures have been detailed in existing surveys [139][45][57]. This section focuses on analysing attack vectors and technical details that are specific to AID systems.
For systematic analysis, the attack vectors are grouped into three categories corresponding to the fundamental principles of cybersecurity: confidentiality, integrity, and availability. Table IV provides this classification and links the attacks identified in Table III with feasible countermeasures summarised in Table V. The following discussion focuses on these attack vectors and outlines potential countermeasures, whereas Section V reviews research contributions on implementing such defences in AID systems.
TABLE III: Documented malicious attacks on AID systems.
Attack Name Reference Source Type Cause of Failure Impact Adversarial Capabilities
RL-based control model attacks Chang et al. [24] (10/2024) Literature (Experimentally Validated) The control system is vulnerable to anomalous inputs. A maximum CGM reading of 400mg/dL caused high rates (over 40%) of hypoglycaemic failure within a few hours; a very slight perturbation (±1mg/dL) can lead to disproportionate patient harm over time High - Read access to the weights of the trained model or estimated weights
AI/ML model attacks Elnawawy et al. [35] (01/2024) Literature (Experimentally Validated) Introducing adversarial CGM data points during inference through a known vulnerability in the Bluetooth communication channel Wrong diagnosis or treatment Medium - Can exploit the Bluetooth communication stack to intercept and manipulate the CGM measurements
Unauthorised remote control Medtronic [84] (01/2023) Real-world (Recall) The settings and insulin delivery of the pump could be potentially controlled through a wireless connection nearby Overdosing or Underdosing Medium - Intercept and manipulate the communication between devices
Process-aware attacks Stergiopoulos et al. [121] (11/2023) Literature (Experimentally Validated) Utilising knowledge of the dynamic process to alter sensory data or the logic inside the controller in an intended but subtle way Overdosing or Underdosing High - Prior knowledge of the components’ implementation mechanism
Replay attacks Walenda et al. [137] (11/2023) Literature (Experimentally Validated) Sending repeating insulin dosages to the insulin pump through a vulnerability in the wireless protocol Replay attacks have the greatest effect when the injected insulin doses are high Low – Intercept the wireless connection and report past packets
DoS attacks Walenda et al. [137] (11/2023) Literature (Experimentally Validated) Overload the system to stop insulin delivery for extended periods DoS attacks are most effective at low, increasing doses, causing longer out-of-range times and greater spikes and dips than replay attacks. Low - Jam the communication channel
Personalised insulin dose manipulation attacks Levy-Loboda et al. [72] (08/2022) Literature (Theoretical) Craft attacks that only threat target patients and are capable of evading extreme overdose/underdose detection Overdosing or Underdosing High – Prior knowledge and manipulation of the insulin dosage
Bias injection attacks Emre Tosunal [126] (06/2022) Literature (Theoretical) A constant bias is injected into the compromised sensor readings Overdosing or Underdosing Medium - Intercept and manipulate the communication between devices
Ripple false data injection attacks Kashyap et al. [60] (05/2021) Literature (Experimentally Validated) Using minimal input perturbations to stealthily change the DNN-predicted insulin dosage More allowed perturbations make attacks easier to find for targeted attacks, with discovery time as short as 60ms. Random attacks can easily induce 10–20 mg/dL output changes across scenarios. High - Read access to the weights of the trained model or estimated weights
Command manipulation attacks Khan et al. [62] (12/2020) Literature (Theoretical) Change the instructions or data that implement the algorithmic steps Overdosing or Underdosing High - Access to the relevant resources; modify a command (i.e., instructions that implement the algorithmic steps)
Unauthorised remote control Cooke et al. [27] (04/2020) Literature (Experimentally Validated) An adversary can remotely control insulin delivery and access unencrypted medical data Overdosing or Underdosing Medium - Intercept and manipulate the communication between devices
Replay attacks Medtronic [85] (08/2018) Real-world (Notification) Involuntary bolus insulin could be delivered to the pump user by copying the wireless radio frequency (RF) signals from remote controllers in proximity Overdosing Low – Intercept the wireless connection and report past packets
CGM attacks Reverberi and Oswald [112] (08/2017) Literature (Experimentally Validated) The Dexcom G4 protocol does not employ cryptographic algorithms Enables the tracking of a user, the jamming of the communication channels and the forging of incorrect sensor readings Low to medium - Software-defined radio and low-cost RF chipsets, reverse engineering
Saturation-based sensor spoofing attacks Park et al. [103] (08/2016) Literature (Experimentally Validated) Make a sensor ignore legitimate inputs by using an additional infrared source to saturate the sensors Inject up to 3.33 times the intended amount of fluid or 0.65 times for a 10-minute period Low – Saturating sensor from an external source
Reverse engineering exploitation Marin et al [81] (03/2016) Literature (Experimentally Validated) Fully reverse-engineered the wireless communication protocol between all the peripherals of the insulin pump system Compromise both the security and privacy of the patient Passive attacks: Low - Radio interception and generation device
@@
Reverse engineering exploitation Li et al [73] (07/2011) Literature (Experimentally Validated) No standard cryptographic mechanisms; the wireless protocol can be fully reverse-engineered, enabling eavesdropping, replay, and command manipulation attacks. Compromise both the security and privacy of the patient Passive attacks: Low - Radio interception and generation device
@@
TABLE IV: Categorised attack vectors and common countermeasures in AID systems.
Security Aspect Attack Vector Description Countermeasures
Confidentiality Eavesdropping
@@
Secure communication protocols;
Proximity communication channels;
Do-it-yourself hacking[66] Patients extract data not provided by the products Authentication
Integrity Replay attacks[137][97] Intercept and retransmit valid data or commands
Bias injection attacks[126] A constant bias was injected into the sensor readings Intrusion detection systems;
Control strategies assess;
Additional sensor for cross-check
False data injection attacks [72][60] Change transmitted data with random or synthesised data
AI/ML model attacks [24][35] Deviate the ML model’s prediction during training or inference
Computational attacks[62] Compromise the computational command or internal data Formal specification
Malicious pump driver [134] Change pump I/O operations
Availability Denial-of-Service (DoS) /Distributed DoS [134] Flooding the target with traffic results in the inability to respond timely Secure communication protocols;